What is the increase in volume when the temperature of 800 mL of air increases from 27°C to 47°C under constant pressure of 1 bar?

Grade:12

2 Answers

Abhinav
6 years ago
160/3 ml is the increase apply the gas equation and substitute values in v1/ v2 = t1/t2 this is only valid when the pressure is kept constant
Prince
5 years ago
T1=27+273=300k
T2=47+273=320k
v1/t1 = v2/ t2
800/300=v2/320
v2=800×320÷300
=2560÷3
=853.3ml
increase in volume=v2-v1=853.3-800=53.3ml